A sample of bikes (2 wheels per bike) is chained
Answer #1

Formula for Ammonium Carbonate (NH4)2CO3

Molar mass of Ammonium carbonate = (14+4)*2 + 12 + 48

=> 36 + 12 + 48

=> 96 gm/mol

moles of ammonium carbonate = 0.702/96 = 0.0073125 moles

Molar mass of FePO4 = 56 + 31 + 64 = 151 gm/mol

Mass of 0.0286 moles = 0.0286 mol * 151 gm/mol = 4.3186 grams

Number of moles of aspirin = 2.32 * 10^(23)/6.023 * 10^(23) = 0.38519 moles

Molar mass of aspirin = 9 * 12 + 8 * 1 + 4 * 16 = 108 + 8 + 64 = 180 gm/mol

Mass of aspirin = 0.38519 mol * 180 gm/mol = 69.334 grams

Molar mass of Diazepam = Mass/number of moles = 15.86/0.05570 = 284.73 gm/mol